# CulturalBench: a Robust, Diverse and Challenging Benchmark on Measuring the

### v1

@@ Abstract
To make large language models (LLMs) more helpful across diverse cultures, it is essential to have effective cultural knowledge benchmarks to measure and track our progress. Effective benchmarks need to be robust, diverse, and challenging. We introduce CulturalBench: a set of 1,227 human-written and human-verified questions for effectively assessing LLMs’ cultural knowledge, covering 45 global regions including the underrepresented ones like Bangladesh, Zimbabwe, and Peru. Questions - each verified by five independent annotators - span 17 diverse topics ranging from food preferences to greeting etiquettes. We evaluate models on two setups: CulturalBench-Easy and CulturalBench-Hard which share the same questions but asked differently. We find that LLMs are sensitive to such difference in setups (e.g., GPT-4o with 27.3% difference). Compared to human performance (92.6% accuracy), CulturalBench-Hard is more challenging for frontier LLMs with the best performing model (GPT-4o) at only 61.5% and the worst (Llama3-8b) at 21.4%. Moreover, we find that LLMs often struggle with tricky questions that have multiple correct answers (e.g., What utensils do the Chinese usually use?), revealing a tendency to converge to a single answer. Our results also indicate that OpenAI GPT-4o substantially outperform other proprietary and open source models in questions related to all but one region (Oceania). Nonetheless, all models consistently underperform on questions related to South America and the Middle East.
@@ 1 Introduction
Uneven cultural representation has been a notorious recurrent limitation of LLMs (Santy et al., 2023; Cao et al., 2023; Arora et al., 2023). Yet, establishing a quality benchmark to effectively gauge LLMs’ nuanced multicultural knowledge remains a formidable challenge (Hershcovich et al., 2022). Effective benchmarks need to be robust, diverse, and challenging. We believe the previous and existing cultural benchmarks may not be satisfactory to be effective. The concrete consequence is that no recent major LLM releases have included cultural evaluation performance in their technical reports (OpenAI et al., 2023; Dubey et al., 2024; Anthropic, 2024). Conventional human-written benchmarks are static and often fail to keep pace with the evolving capabilities of LLMs (Yang et al., 2023). Alternatively, existing auto-generated benchmarks cannot reflect the real struggles of models and the true concerns of users on multicultural knowledge. They often rely on web resources e.g., Wikipedia (Naous et al., 2023; Fung et al., 2024), and LLMs’ responses on established human surveys e.g., World Value Survey (Durmus et al., 2023b; Li et al., 2024). Those benchmarks could be less effective since the scraped web sources have been used directly on training and the surveys have limited cultural concepts. The latest synthetic data benchmark approach (Rao et al., 2024; Fung et al., 2024), despite their scalability, risk propagating existing data distribution bias in models that they are meant to measure (Liu et al., 2024).
@@
Drawing insights from recent red-teaming approaches on LLMs’ safety (Ganguli et al., 2022) and interactive model evaluation and data collection efforts (Kiela et al., 2021; Chiang et al., 2024), we develop CulturalTeaming, an AI-assisted interactive red-teaming data collection and validation pipeline. CulturalTeaming aims to construct a robust, diverse and challenging benchmark. The pipeline consists of three parts as shown in Fig. 1 – (1) Red-teaming data collection (2) Human Quality Check (3) Filtering. The goal of the red-teaming platform is to guide and encourage humans in iteratively create challenging questions for models. Specifically, humans provide diverse cultural scenarios based on their daily observations and unique cultural knowledge. The AI helper provides writing assistance to alleviate the burden of formulating questions.
@@
We introduce our collected CulturalBench with 1,227 high-quality questions, each of which has been verified by five independent annotators. These questions span 45 global regions including less represented ones such as Bangladesh in South Asia, Zimbabwe in Africa, and Peru in South America, with details in Fig. 6. They are diverse with 17 cultural topics identified in Fig. 3, that reflect a broad spectrum of cultural elements in different countries/regions e.g., food, language/communication, visiting etiquette, and celebrations.
@@
To capture the cultural diversity in each region, our CulturalBench contains two types of questions: (i) Single-mode: one correct answer and (ii) Multiple-mode: multiple correct answers, as shown in Fig. 1. During the human quality check, we allow annotators to respond to each question in a multi-label format, recognizing that multiple valid answers can coexist for some questions (Boratko et al., 2020). For instance, for a question of “what utensil do Chinese people usually use everyday?”, the most likely answer is “chopsticks” (which is a common utensil for eating Chinese food). However, other answers such as “spoon” may also reflect the reality of the Chinese population, depending on the specific foods being served.We have strict criteria on filtering out questions with no answer having majority vote (i.e., $\geq$ 4 out of 5 annotators), ensuring our CulturalBench is robust and captures accurate cultural representations.
Figure 1: Overview of AI-assisted red-teaming data collection and validation to construct CulturalBench.
@@
There are two evaluation setups on our CulturalBench– (1) CulturalBench-Easy, which evaluate the model on multiple choice questions; (2) CulturalBench-Hard, which converts the multiple choice question into binary questions (True/False) for each of the four options as shown in Fig. 1. After collecting data, we first designed and constructed our CulturalBench-Easy, directly using the 1,227 standardized questions with four options. Although there are performance differences (28%) between the worst and best-performing models, the best-performing model achieves 88%, which only slightly lags behind the human baseline (92.4%). Inspired by the recent studies on binary setting to accurately test models’ reasoning capabilities (Kadavath et al., 2022; Zhang et al., 2024), we construct our CulturalBench-Hard by converting the 1,227 multiple-choice questions to 4,908 binary questions (four per original question). We test 30 models from different families (e.g., OpenAI GPT, Llama, Qwen) across different model sizes (e.g., 8b, 70b, and 405b). We found this setup to be much more challenging for LLMs with the best performing model at only 61.5% accuracy and the worst at 21.4%, compared to a human performance of 92.6%.
@@
Looking to understand why models perform drastically different on CulturalBench-Easy and -Hard, we wondered if models can simply guess the most likely option under multiple-choice format found in the CulturalBench-Easy setup. We designed an experiment that shows that models can get 40% accuracy (substantially above random chance of 25%) by simply choosing the option that has greatest embedding similar to the name of the culture (without seeing the question). This shows the potential limitation of assessing models’ capabilities under the multiple-choice setting in CulturalBench-Easy since they could rely on such heuristics without needing to demonstrate cultural understanding. In contrast, CulturalBench-Hard can more effectively assess the cultural knowledge of models, because such heuristics cannot be easily applied to game evaluation.
@@
Moreover, our evaluation on different question types shows that even the best models struggle with questions that have multiple correct answers, revealing a tendency to LLMs to over-converge on a single option. This is evident by a significant drop (-19.8%) in accuracy on questions with multiple correct answers, as compared with questions with a sole correct answer. Through our analysis of questions relating to various sub-continents in CulturalBench-hard, we find that models perform well on questions relating to regions (e.g., North America and South Asia) that are highly represented in web-source data (e.g., United States, as part of North America) and large-scale human annotation sources (e.g., India in South Asia). However, models underperform on questions relating to less well-represented regions such as Eastern Europe. This observation holds even for models developed by providers based outside of the United States (e.g. Alibaba Qwen and Mistral), which might possibly be attributed to the availability of the data used in various stages of training. Overall, OpenAI’s GPT-4o outperforms other proprietary providers and open source model builders uniformly across all but one region (Oceania). With CulturalBench and our analysis on various models, we provide an effective benchmark for testing the cultural knowledge of various LLMs, with the hopes of encouraging model developers to easily perform cultural evaluations in the journey to develop more culturally-sensitive LLMs.

### v2

# CulturalBench: A Robust, Diverse, and Challenging Cultural Benchmark by Human-AI CulturalTeaming

@@ Abstract
Robust, diverse, and challenging cultural knowledge benchmarks are essential for measuring our progress towards making LMs that are helpful across diverse cultures. We introduce CulturalBench: a set of 1,696 human-written and human-verified questions to assess LMs’ cultural knowledge, covering 45 global regions including underrepresented ones like Bangladesh, Zimbabwe, and Peru. Questions are each verified by five independent annotators and span 17 diverse topics ranging from food preferences to greeting etiquette. We construct CulturalBench using methods inspired by Human-AI Red-Teaming. Compared to human performance (92.4% accuracy), the hard version of CulturalBench is challenging even for the best-performing frontier LMs, ranging from 28.7% to 61.5% in accuracy. We find that LMs often struggle with tricky questions that have multiple correct answers (e.g., What utensils do the Chinese usually use?), revealing a tendency to overfit to a single answer. Our results indicate that GPT-4o substantially outperform other models across cultures, besting local providers (e.g., Mistral on European culture and DeepSeek on Chinese culture). Across the board, models under-perform on questions related to North Africa, South America and Middle East. [^1] We publicly release the CulturalBench data with Creative Commons Attribution 4.0 license.
CulturalBench: A Robust, Diverse, and Challenging
Cultural Benchmark by Human-AI CulturalTeaming
Yu Ying Chiu1 Liwei Jiang1 Bill Yuchen Lin1 Chan Young Park1 Shuyue Stella Li1 Sahithya Ravi2,3 Mehar Bhatia4,5 Maria Antoniak6 Yulia Tsvetkov1 Vered Shwartz2,3 Yejin Choi7 1 University of Washington 2 University of British Columbia 3 Vector Institute for AI 4 McGill University 5 Mila 6 University of Copenhagen 7 Stanford University kellycyy@uw.edu lwjiang@cs.washington.edu [Uncaptioned image] https://hf.co/spaces/kellycyy/CulturalBench
@@ 1 Introduction
Figure 1: The human-AI collaborative data collection pipeline of CulturalBench.
@@
Uneven cultural representation has been a notorious recurrent limitation of LMs (Santy et al., 2023; Cao et al., 2023; Arora et al., 2023). Yet, establishing a quality benchmark to effectively gauge LMs’ nuanced multicultural knowledge remains a formidable challenge (Hershcovich et al., 2022). Effective benchmarks need to be robust, diverse, and challenging. Conventional human-written benchmarks are static and often fail to keep pace with the evolving capabilities of LMs (Yang et al., 2023). Alternatively, existing auto-generated benchmarks cannot sufficiently challenge existing models and reflect aspects of multicultural knowledge that users are concerned about. Instead, they often rely on web resources e.g., Wikipedia (Naous et al., 2023; Fung et al., 2024), and LMs’ responses on established human surveys e.g., World Value Survey (Durmus et al., 2023b; Li et al., 2024). Those benchmarks could be less effective since scraped web sources have been used directly on training and the surveys have limited cultural concepts. Despite their scalability, the latest synthetic data benchmark approaches (Rao et al., 2024; Fung et al., 2024) risk propagating existing data distribution bias in models that they are meant to measure (Liu et al., 2024).
Contribution 1: We develop CulturalTeaming, a collaborative human-AI red-teaming data collection pipeline.
@@
We draw insights from recent red-teaming approaches on LMs’ safety (Ganguli et al., 2022) and interactive model evaluation and data collection efforts (Kiela et al., 2021; Chiang et al., 2024). The pipeline consists of three parts as shown in Fig. 1 – (1) Red-teaming Data Collection (2) Human Quality Check (3) Filtering. The goal of the red-teaming platform is to guide and encourage humans to iteratively propose challenging questions for models. Specifically, humans provide diverse cultural scenarios based on their daily observations and unique cultural knowledge. The AI helper provides writing assistance to alleviate the burden of formulating questions.
Contribution 2: We introduce CulturalBench containing 1,696 high-quality, challenging and diverse questions with full verification by human annotators.
@@
Each question is verified by five independent annotators. These questions span 45 global regions including less represented ones such as Bangladesh in South Asia, Zimbabwe in Africa, and Peru in South America, with details in Fig. 4 and Appendix C. These questions cover 17 cultural topics identified in Fig. 6, reflecting a broad spectrum of cultural elements, such as food, social etiquette and celebrations.
@@
CulturalBench contains two question types: (i) Single-mode: one correct answer and (ii) Multiple-mode: multiple correct answers, as shown in Fig. 5. During the human quality check, we allow annotators to respond to each question in a multi-label format, recognizing that multiple valid answers can coexist for some questions (Boratko et al., 2020). For instance, for a question of “what utensil do Chinese people usually use everyday?”, the most likely answer is “chopsticks” (which is a common utensil for eating Chinese food). However, other answers such as “spoon” may also reflect the reality of the Chinese population, depending on the specific foods being served. We have strict criteria on filtering out questions with no answer having majority vote (i.e., $\geq$ 4 out of 5 annotators), ensuring our CulturalBench is robust and captures accurate cultural representations.
Contribution 3: We reveal uneven performance of models on cultural knowledge across question types and regions.
@@
There are two evaluation setups for CulturalBench: (1) CulturalBench-Easy, which evaluates the model on multiple choice questions; (2) CulturalBench-Hard, which converts each multiple choice question into a multi-label question with binary choice (True/False) for each of the four options as shown in Fig. 5. After collecting data, we first designed and constructed our CulturalBench-Easy, directly using the 1,696 standardized questions with four options. Although there are performance differences between the worst and best-performing models, the best-performing model achieves 89.6%, which only slightly lags behind the human baseline (92.4%). Inspired by the recent studies on binary setting to accurately test models’ reasoning capabilities (Kadavath et al., 2022; Zhang et al., 2024), we construct our CulturalBench-Hard by converting the 1,696 multiple-choice questions to 6,784 binary questions (four per original question). We test 29 models from different families (e.g., GPT, Llama, Cohere, Deepseek) across different model sizes (e.g., 8b, 70b, and 405b). We found this setup to be much more challenging for LMs with the best performing model (OpenAI o1) at only 61.4% accuracy and the worst at 28.7% (Aya-8b), compared to human performance of 92.4%.
@@
Looking to understand why models perform so differently on CulturalBench-Easy and -Hard, we analyze if models can simply guess the most likely option under multiple-choice format in the CulturalBench-Easy. We show that a trivial baseline can get 40% accuracy (substantially above the random chance of 25%) by choosing the option that has the greatest embedding similarity with the name of the culture (e.g., chopsticks/spoon with China), without at all needing the question. This shows the potential limitation of assessing models’ capabilities under the multiple-choice setting in CulturalBench-Easy since they could rely on such heuristics without needing to demonstrate cultural understanding. In contrast, CulturalBench-Hard can more effectively assess the cultural knowledge of models, because such heuristics cannot be easily applied to game evaluation.
@@
Moreover, our evaluation on different question types shows that even the best models struggle with questions that have multiple correct answers, revealing a tendency for LMs to over-converge on a single option. This is evident by a significant drop (-28.7%) in accuracy on questions with multiple correct answers, as compared with questions with a single correct answer. Through our analysis of questions relating to various sub-continents in CulturalBench-hard, we find that models perform well on questions relating to regions (e.g., North America and South Asia) that are highly represented in web-source data (e.g., United States, as part of North America) and large-scale human annotation sources (e.g., India in South Asia). However, models underperform on questions relating to less well-represented regions such as East Europe. Surprisingly, this observation holds even for models developed by providers based outside of the United States for the questions relating to certain regions (e.g. Mistral for West Europe, and Qwen for East Asia), which might possibly be attributed to the availability of the data used in various stages of training. Overall, OpenAI GPT-4o outperforms other proprietary providers and open-source model builders uniformly across all regions.
With CulturalBench, we take the first step toward providing an effective and high-quality benchmark for testing the cultural knowledge of various LMs. We hope to encourage other researchers to build benchmarks that integrate human-AI efforts inspired by our human-in-the-loop red-teaming CulturalTeaming in the journey toward more culturally sensitive LMs.
2 Data collection pipeline
Our data collection pipeline consists of three steps, as illustrated in Fig. 1: (1) Data collection via human-AI red-teaming (2) Human quality check on full data (3) Filtering with majority vote. Such a multi-step process enables us to collect robust data for CulturalBench.
2.1 Step 1: Data collection via human-AI red-teaming
Question Formulation.
Human annotators are instructed to brainstorm culturally relevant scenarios based on their personal experiences of their cultures (e.g., Singaporeans use tissue packet to reserve seats). A step-by-step guideline with detailed examples is provided to inspire them, as shown in Appendix E. The AI helper bot then transforms the scenario into a structured question with four options, which the annotators can review and edit.
Question Verification & Revision.
Human annotators can use the formulated question as basis to challenge the AI verifier in our interactive platform. The platform provides further assistance in revising the questions to make it more challenging by offering various revision strategies along with drafted examples (e.g., “Negate the Question”), as shown in Appendix E.
Internal Filtering.
After collecting over 3,600 questions, the researchers carefully reviewed and removed those that are not relevant to any regions (e.g., Bangladesh, Peru and Hong Kong), resulting in a filtered set of over 3,000 cultural questions.
2.2 Step 2: Human Quality Check
Recruitment Criteria.
We collected questions at the country/regional level, pairing each question with a specific region. To ensure culturally attuned and thorough verification, we recruited five annotators for each region through the Prolific platform  [^2] https://www.prolific.com. We set two main criteria to ensure that the recruited annotators have a deep understanding of the culture of the targeted country or region – (1) Nationality (2) Primary residence before age 18. For certain cultures (e.g. the United States, the United Kingdom), when our collected questions have targeted specific groups in the country/region, we try to have more fine-grain selection to find the target group such as ethnicity (e.g. African American), and place of residence (e.g., Wales). See more detail in Sec. Further fine-grained culture classification..
Multiple Selection Settings.
To better reflect the true representation of each cultural question, we allow annotators to select multiple answers on our questions with four options. As a result, some questions may have more than one majority-vote answer. We also give them the choice to select “e. the question has no correct option (or is otherwise unanswerable)” or “f. they have no knowledge to answer this question”, to avoid them guessing. See more details in Appendix Fig. 13 and 14. This approach also helps test models’ mode-seeking behavior, examining whether they rely solely on cultural stereotypes (i.e., modes) without considering broader cultural diversity.
2.3 Step 3: Filtering by Majority Vote & Constructing Benchmarks
Majority Vote Criteria.
To build a robust benchmark that captures the accurate representation of cultural knowledge, we set the majority-vote threshold to be $\geq 4$ out of 5 annotators. During human validation, we first filtered out questions without majority votes, resulting in a final set of 1,696 questions. Subsequently, we further processed the remaining questions. To construct our CulturalBench in two setups (CulturalBench-Easy: Multiple-choice, CulturalBench-Hard: True/False), we processed the questions differently depending on the number of majority-vote answers they contain.
(1) Single-Mode Questions (Only one majority-vote answer).
  • CulturalBench-Easy. We directly keep the original question with four options.
  • CulturalBench-Hard. We transform the question with four options into four binary questions. For instance, the question drafted (e.g., “What do Singaporeans …? A. Tissue … D. Book”) will form four binary questions (e.g. “Is this answer true or false for this question? Answer True or False only. Question: What do Singaporeans …? Answer: Tissue.”)
(2) Multi-Mode Questions (More than one majority-vote answer).
  • CulturalBench-Easy. We reframe the question to allow selecting multiple statements. The four drafted options (e.g., “A. Tissue”) become the four statements in questions (e.g., “(i) Tissue”). To ensure the models know the possibility of questions containing multiple correct labels, we add the instruction on question directly with (“Select the options with all applicable statements”) with some options being a composite statement (e.g., “A. (i) Tissue and (iv) Book”).
  • CulturalBench-Hard. We follow the same construction approach as single-mode questions.
Figure 2: Models performance on CulturalBench-Hard with random baseline at 6.25% and human performance at 92.4%.
3 CulturalBench Description and Discovered Topics
3.1 Descriptive Statistics
Our benchmark covers a wide range of global regions, spanning 45 countries and regions, including underrepresented regions such as Bangladesh, Zimbabwe, and Peru. A detailed breakdown of regional distribution can be found in Appendix C while example questions by topic are in Appendix B.
CulturalBench-Easy.
It contains 1,696 multi-choice questions, each with four options. The gold label is the correct option (A, B, C or D).
  • Single-mode. In Fig. 5, a question of “What do Singaporeans usually use to reserve seats?” with options of “A. Tissue … D. Book”.
  • Multi-mode. In Fig. 5, a question of “What do Singaporeans…? Selecting the option with all applicable statements. i) Tissue … iv) Book” with options of “A. (i), … D. (i), (iv)”.
CulturalBench-Hard.
This set contains 1,696 ×4$\times$ 4=× 4 = 6,784 True/False judgement questions. For evaluation, models need to answer all four binary questions correctly to count as knowing the cultural knowledge of one question. More detail can be found in Sec. 4.
  • Single-mode. “Is this answer true or false for this question? Question: What do Singaporeans usually use to reserve seats? Answer: Tissue.”, as shown in Fig. 1. There is only one “True” for the four transformed binary questions.
  • Multi-mode. There are more than one “True” for the four transformed binary questions.
3.2 Diverse Topics Discovered Across Cultures
Most existing cultural benchmarks have predefined topics to collect data on, typically on universal topics such as dining (Adilazuarda et al., 2024). However, this approach can overlook cultural elements unique to specific regions. To capture a broader spectrum of cultural topics, we adopted a discovery-based approach by encouraging human annotators to brainstorm cultural concepts from their personal experiences. The detailed instruction for annotators can be found from Fig. 8 to Fig. 12 in Appendix E. CulturalBench spans a diverse range of cultural elements with 17 topics under three overarching categories (Daily life, Social Etiquette, and Wider Society), as shown in Fig. 6. Daily life relates to the everyday experiences of people e.g., Workplace. Social Etiquette means the acceptable norms in society e.g., Greeting. Wider Society includes special elements for broader spectrum of cultural topics e.g., Celebrations. We classified questions into topics by prompting GPT-4o. The classification prompt and the topic detailed definitions are in Appendix B.
To collect diverse data for each culture, we allow each annotator to create at most 3-7 questions, depending on the availability of annotators for each region. Notably, in curating CulturalBench, we observed that people from different regions focused on distinct topics. For instance, annotators from Italy provided more questions related to Food, with 38.9% (14) questions of total. In contrast, participants from Israel focused more on Religion, contributing 23.8% (10) questions of total. Our discovery-based approach allow us to capture diverse cultural elements from people in different regions without being limited by a predefined set of topics.
4 Experiments: Evaluation of LMs on CulturalBench
We evaluate 29 current LMs in a zero-shot setting on CulturalBench in two setups: (1) CulturalBench-Easy: Multiple choice; (2) CulturalBench-Hard: True/False, as shown in Appendix Tables 10 and 11. We prompted the models to ensure they follow the output format to allow fair comparison. The detailed prompt is in Appendix D. To avoid exposing the correct answers to models for fair comparison, our annotation platform, which involves using OpenAI APIs, disallowed the collected data to be used for training.
CulturalBench-Easy.
We evaluate model performance by measuring accuracy, specifically whether the model correctly identifies the label for each multiple-choice question. A random baseline can achieve 25%.
Figure 3: Analysis on question type (Left) and time version (Right). For question type, we demonstrate models struggle at answering questions with multi-modes (more than one correct answers). For time version, we show the improved performance of models across time, but some are approaching a performance ceiling.
CulturalBench-Hard.
We evaluate model performance based on the proportion of tasks in which the model can get all four options predicted correctly. For each task, an LM has to make four binary judgements (True/False) from the transformation of four options in each multiple choice question. To demonstrate robust cultural knowledge, we believe the LM has to accurately which option(s) are False as well as which option(s) are True. A random baseline can achieve $0.5^{4}$ = 6.25%.
4.1 Comparing LMs on two benchmarks across model family and size
Due to space constraint, we show the performance of 23 representative models across model families and sizes on CulturalBench-Hard in Fig. 2. The corresponding Fig. for CulturalBench-Easy is in Fig. 7.
Models shows a huge performance gap with humans on CulturalBench-Hard.
As shown in Fig. 2, this setup is significantly more challenging for current LMs, with accuracy ranging from 28.7% for Cohere Aya-8b to 61.4% for OpenAI o1. These scores are considerably lower compared to the human baseline of 92.4%, highlighting the difficulty of the task even for the most advanced models.
Models performance improves as model size increases.
In Fig. 2, we present the performance of models from six different families, such as GPT, Llama, and Qwen. Overall, the results demonstrate a trend of improved performance as model size increases. For example, within the OpenAI model family, the models show a clear progression in accuracy: GPT-3.5 Turbo achieves 34.5%, GPT-4o attains 60.4% and o1 attains 61.5%. Similarly, in Fig 2, in the Llama-3.1 family, 8B achieves 36.0%, 70B has 54.6%, and 405B has 51.93%. The 70b model has a much higher score than the 8b model, albeit the inconsistent pattern in Llama-3.1 405B. This pattern is consistent across most of the model families, indicating that larger models generally have better cultural knowledge.
All models tested have a huge performance difference between the two setups of CulturalBench.
For instance, the best-performing model, o1, achieves 89.6% accuracy on CulturalBench-Easy but drops to 61.4% accuracy on CulturalBench-Hard, whereas the human baselines are both 92.4% and the random baselines are 25% and 6.25% respectively. We hypothesize that the models can guess the most possible answer on CulturalBench-Easy under the multiple-choice setting. To investigate this hypothesis, we compute the embedding for the country name and separately for each option using OpenAI text-embedding-3-small.
By using a simple heuristic of choosing the option with highest cosine similarity with the country name (e.g. Bangladesh), we attain 40.4% accuracy. This is intriguing as it is substantially above the random baseline for multiple-choice setup in CulturalBench-Easy (25%), without needing considering the question at all. We find that the cosine similarity difference between the correct option and the country name is significantly higher than the difference between options average and the country (0.166 vs. 0.145; Kruskal-Wallis $p$$\leq$0.01). This shows the possibility of models guessing based on one (out of many possible) heuristics in multiple-choice setup without understanding (or even knowing) the question. This stresses the importance on using the binary (True/False) for each of the four options per question in CulturalBench-Hard to accurately assess cultural knowledge of LMs.
Figure 4: Models performance by different providers on various geographic regions. We further compare seven representative models (GPT-4o, Claude Sonnet, Mistral Large, Cohere Aya 32b, Llama-3.3-70b, DeepSeek V3 and Qwen-2.5 72b) from different model families. We avoid plotting OpenAI o series and DeepSeek R1 as they spend substantial reasoning tokens, which might be unfair comparison.
4.2 Investigating effects of question type and time version of models
LMs show distinct gaps between question types, unlike humans.
We evaluate the performance based on question types – (1) Single-mode (N=1554) and (2) Multi-mode (N=142). The first type refers to the questions with only one correct answer while the second type includes questions with multiple correct answers, as explained in Section 2.3. All the correct answers are majority-voted. In Fig. 3 (Left), the average across all models shown in Fig. 2 is 49.6% on Single-mode questions and 20.9% on Multi-mode questions, revealing a significant gap of 28.7% between the two. Similarly, the best model (o1) exhibits a 45.5% performance difference between these question types. In contrast, human baselines show only a 6.1% difference, indicating that humans handle cultural nuances (where one question could have more than one correct answers) more effectively than models. This discrepancy suggests that models struggle to account for cultural nuances due to their mode-seeking tendencies, supporting a similar observation by Tajwar et al. (2024).
Model providers improve on cultural knowledge of models across time, but are approaching a performance ceiling.
In Fig. 3 (Right), we evaluate three popular models familiy (GPT, Llama, Deepseek) across released time. Overall, all model families demonstrate an increasing trend in performance across time. However, most model families show gradual improvements for their newer versions. For instance, the Llama family shows only a slight improvement from Llama-3.1-70b to Llama-3.3-70b, compared to the improvement between Llama-3-70b and Llama-3.1-70b.
4.3 Studying different providers’ LMs on questions from different regions
We include detailed performance of models across different family and sizes (shown in Fig 4) to understand how well different models performance in questions relating to different geographic regions at a sub-continent level.
Models perform better in questions relating to North America, North Europe and South Asia.
From Fig. 4, it is evident that models achieve higher performance averages in regions like North America (57.9%), North Europe (51.8%) and South Asia (51.5%). We hypothesize that the higher performance in these regions can be attributed to several factors including their representation on web-data used for model training (Longpre et al., 2023) and the proportion of annotators recruited from these regions by LM providers to curate post-training alignment data. For instance, many annotators are known to be recruited from India as they have good English ability and costs substantially less than their counterparts in the United States (Lohchab and Roy, 2024).
Models score lower in questions relating to South America, East Europe, and the Middle East.
Models exhibit lower performances on average in regions like South America (41.5%), East Europe (41.5%), and Middle East/West Asia (37.8%). These disparities possibly suggest insufficient representation of cultural knowledge from these regions in training data at various stages.
GPT-4o leads in most regions, followed by Llama-3.3 70b and Claude-3.5 Sonnet.
GPT-4o consistently ranks highest across most regions among all tested models. Llama-3.3 70b shows strength in regions where cultural knowledge is traditionally less represented, such as South Africa, while Claude-3.5 Sonnet performs particularly well in other regions e.g., West Europe, West Africa.
Chinese and European Model Providers are not stronger in cultural knowledge relating in their respective regions.
Despite claims of specialization in local languages, Qwen-25-72b, Deepseek V3 and Mistral Large do not outperform other models in their respective regions in terms of cultural knowledge. For example, Qwen-2-72b and Deepseek V3 score 58.8% and 61.3% on East Asia, while GPT-4o achieves 64.6%. Similarly, Mistral Large underperforms in West Europe (44.6%) compared to GPT-4o (56.2%). These results suggest that model providers based out of specific regions do not necessarily have advantages in cultural knowledge of their regions.